WAEC Further Mathematics 2009 Theory question 105

Theory question 105

Probability

The probabilities of Rotey obtaining the highest mark in Mathematics, Physics and Biology tests are 0.9, 0.75 and 0.8 respectively. Calculate the probability of getting the highest marks in at least two subjects.

Show the model answer
Model answer

All three: 0.9 × 0.75 × 0.8 = 0.54
M, P not B: 0.9 × 0.75 × 0.2 = 0.135
M, B not P: 0.9 × 0.25 × 0.8 = 0.18
P, B not M: 0.1 × 0.75 × 0.8 = 0.06
At least two = 0.915

Explanation

'At least two' means exactly two or all three.